A student is trying to determine the concentration of an acetic acid (HC2H3O2) solution. They place 5.00 mL of it in a flask and titrate it with a 0.150 M NaOH solution. At the endpoint of the titration, they find that they have used 19.27 mL of the NaOH solution.

Based on this information answer the following questions.

a) How many moles of NaOH are used to reach the endpoint? ____moles NaOH

b) How many moles of acetic acid were neutralized by the NaOH added? _____moles HC2H3O2

c) What is the molar concentration of the original acetic acid solution? ________M HC2H3O2
Answer #1

a) The moles of NaOH are calculated:

n NaOH = M * V = 0.15 M * 0.01927 L = 2.89E-3 mol

b) The moles of acetic acid are the same as those of NaOH:

n Acid = 2.89E-3 mol

c) Calculate the molar concentration of the acid:

[Acid] = n / V = ​​2.89E-3 mol / 0.005 L = 0.578 M

If you liked the answer, please rate it positively, you would help me a lot, thanks.